OpenStudy (anonymous):

bob deposits $3,632 in a bank account that pays 3.4% simple interest. if she doesnt change the principal, what is her balance after 5 years? a. $617.44 b. $3,755 c. $4,249.44 d. $12,348.80 how would you work out the problem and what would your steps be?

OpenStudy (dumbcow):

for simple interest problems B = P(1+i*t) where i is int rate and t is time so B = 3632(1+(0.034)*5)

OpenStudy (dumbcow):

for simple interest problems B = P(1+i*t) where i is int rate and t is time so B = 3632(1+(0.034)*5)